# Topologic Attention Networks: Attending to Direct and Indirect Neighbors through Gaussian Belief Propagation

### v1

@@ 3.6 Addressing the Unbounded Memory
Although Neural GaBP guarantees convergence under walk-summability, the number of iterations required to reach convergence is unknown. This poses a challenge for training, as it would require storing an unknown number of iteration messages for each layer. To address this, we adopt an implicit differentiation strategy that computes exact gradients at convergence without unrolling the iterative inference process. At convergence, the solution satisfies $J\mu=h$. Differentiating with respect to the model parameters $\theta$ gives the corresponding forward and backward relations:
@@
This relation allows gradients to be computed directly at the fixed point using the same GaBP solver as in the forward pass, eliminating the need to store intermediate messages. This reduces memory complexity from $O(TE)$ to $O(E+N)$, trading additional compute for substantial memory savings.
@@ 4 Experimental Setup
We evaluate the performance of learned and fixed topologies for TANs across six benchmark datasets with cyclic graph structures under both homogeneous and heterogeneous learning settings. The Cora, Citeseer, and PubMed citation datasets serve as representative homogeneous benchmarks, while the Texas, Wisconsin, and Cornell datasets serve as representative heterogenous benchmarks. These six datasets each represent a single, static graph and are used for transductive node classification.
@@
We compare our method against foundational and representative models from the other two branches of dynamic neighborhood aggregation, along with standard local aggregation methods. Specifically, we include the Graph Convolutional Network (GCN) and Graph Attention Network (GAT) as representative local aggregation models. For methods that perform global or continuous neighborhood aggregation, we evaluate against the Structure-Aware Transformer (SAT) [structure_aware_transformer], an adapted Graphormer designed for node classification rather than graph classification, as well as the Graph Neural Differential Equation (GDE) and Continuous Graph Neural Network (CGNN).
@@
To ensure a fair comparison across all models, we use a consistent architecture with the hidden dimensions fixed at 64. Each method is implemented with two layers of its respective aggregation mechanism. For attention-based architectures, each layer employs eight attention heads in the first layer and a single head in the second layer. Each first layer head outputs eight features, while the second layer head outputs all 64 features. This setup controls for model capacity so the differences primarily reflect the behavior of each aggregation mechanism rather than variations in network depth or parameter count.
@@
All models are trained using the Adam optimizer with weight decay $5\cdot 10^{-4}$, early stopping with a patience of 100 epochs and a dropout of $0.6$, except the fixed Laplacian which used a patience of 200 epochs. The learning rate for the TAN and SAT models are $1\cdot 10^{-3}$ while the learning rate for the other models are $1\cdot 10^{-2}$. The tolerance for the GaBP in the Topologic Attention Layers is set to $1\cdot 10^{-6}$ and a maximum iteration count was set to 1000. To ensure stable convergence on graphs with loops, we apply message damping with a coefficient of $\lambda=0.5$.
@@
For each dataset, we adopt random node splits preserving the original train/validation/test ratios but do not enforce class balance. This setup better reflects realistic scenarios where labeled nodes are unevenly distributed across classes, providing a more robust measure of generalization [shchur2019pitfalls]. Each experiment is repeated across 30 random runs, each with independently sampled splits. We report the mean $\pm$ standard deviation of test accuracy.
@@
For TAN, we ensure the graph is undirected and self loops are removed since the diagonal entries of the precision matrix $J$ already encode each node’s self-precision. For the other models, we ensure self loops.
@@ 5 Results
Table 1 reports the performance across all benchmark datasets. Overall, TAN achieves improved generalization on all measured baselines.
@@
Within TAN, we implemented six variants: a learnable and a fixed versions of each precision matrix. The fixed Laplacian construction, although walk-summable, was the only variant that did not converge within the maximum iterations. This occurs because normalized Laplacian topologies place much of their spectrum very close to the walk-summability boundary, which results in extremely slow GaBP contraction even after symmetric normalization and scaling. As a result, both the forward graph linear transformation and the backward implicit-differentiation were computed using partially converged solutions. This produced noisy graph linear transformations and gradients which made early training difficult; however once the model began to form stable representations, the fixed Laplacian variant was the most general matrix construction tested with state-of-the-art accuracy in five out of six datasets. This suggests that exact GaBP convergence is not always necessary for effective learning.
For the other matrix constructions, the learned precision matrices always converged under the GaBP solver, requiring no more than 100 iterations to reach the tolerance threshold. The fixed Diagonally Dominant and fixed Pairwise Normal constructions also converged consistently, although they typically required more iterations than their learned counterparts, reaching at most about 150 iterations.
@@ 6 Conclusion
This work introduced TANs, which implement topologic attention by learning how information should flow probabilistically through the graph’s topology. TANs achieve this by using Gaussian Belief Propagation to exchange messages between nodes, allowing each node to integrate both direct and indirect influences. Across both homophilic and heterophilic benchmarks, TANs consistently surpassed the measured baselines.
A particularly notable finding was that the fixed Laplacian variant achieved high accuracy even when GaBP did not fully converge. This result suggests that partially converged GaBP solutions may introduce useful inductive biases or implicit regularization effects. Further investigation into these approximate solutions may reveal new opportunities for designing efficient inference procedures that maintain strong generalization performance.
We are excited to see the future development of TANs. One important challenge for wider adoption is improving the efficiency of GaBP on arbitrary graphs. Because GaBP is an iterative algorithm, incorporating multigrid techniques [multigrid] or other hierarchical solvers may lead to significantly faster convergence and better scalability, which would enable TANs to operate effectively on large or streaming graph datasets.
@@ Appendix A Appendix
A.1 Psuedocode for Gaussian Belief Propagation
We provide the psuedocode for GaBP as a convenient understanding for the math presented in the background section.
@@ A.1 Psuedocode for Gaussian Belief Propagation
8:  $\pi_{i\to j}\leftarrow(1-\lambda)\,\pi_{i\to j}+\lambda\,\pi^{\text{new}}_{i\to j}$
9:  $\eta_{i\to j}\leftarrow(1-\lambda)\,\eta_{i\to j}+\lambda\,\eta^{\text{new}}_{i\to j}$
10:  $\Delta=\max_{(i,j)}(|\Delta\pi_{i\to j}|,|\Delta\eta_{i\to j}|)$
11:end while
12:Output: $\mu_{i}=\eta_{i}/\pi_{i}$
A.2 Evolution of GaBP Iterations During Training
Since TANs rely on GaBP to perform the graph linear transformation at each layer. We wanted to analyze how the number of GaBP iterations evolves over the course of training for the learned precision matrices. We visualize (Figures the convergence behavior along with the accuracy on Wisconsin dataset the three learned precision matrices. We chose the Wisconsin dataset to analyze because the learned matrices achieved similar accuracy on the dataset. We ran the models for 500 epochs. We also visualize the diagonally dominant matrix for the Pubmed dataset. Interestingly, across the analyzed datasets, the iterations decrease as training progresses. Moreover, the dataset does not appear to overfit to the dataset.
Figure 4: Training Over Time of the Diagonally Dominant Precision Matrix on the Wisconsin Dataset.
Figure 5: Training Over Time of the Laplacian Precision Matrix on the Wisconsin Dataset.
Figure 6: Training Over Time of the Pairwise Normal Matrix on the Wisconsin Dataset.
Figure 7: Training Over Time of the Diagonally Dominant Precision Matrix on the Pubmed Dataset.
@@ A.3 Analysis of the Biases of Matrix Constructions
To understand how each precision matrix construction shapes the topology, we visualize the correlation matrix (Figures 811) on the Wisconsin dataset. For the learned precision matrices, we visualize the first-layer. These correlation matrices are computed from the inverse precision matrix and describe how strongly two nodes would co-vary under the learned topology in the absence of any observation vector $h$. They therefore characterize the influence structure implied by $J$, meaning the potential pathways and relative strengths along which GaBP is capable of transmitting information, regardless of the actual evidence supplied during inference. Red values indicate positive correlations and blue values indicate negative correlations, with intensity proportional to magnitude. Nodes are ordered by the second eigenvector to highlight community structure, and the accompanying adjacency matrices (bottom panels), ordered in the same way, show the corresponding learned connectivity patterns.
@@
The Laplacian-based precision matrix produces a pronounced block structure resembling a plaid pattern of alternating positive and negative correlations. Nodes form compact clusters with internal correlations, while inter-cluster relationships often alternate in sign, indicating that the model captures both homophilic and heterophilic dependencies. The dense regions along the diagonal of the reordered adjacency matrix confirm the emergence of localized communities. Because the Laplacian construction constrains the eigenvalues of the precision matrix to lie within $(0,1)$, each node contributes with comparable influence to the overall propagation. Most of the learning therefore occurs through topologic deformation, where the model folds the underlying space before blending node representations. As a result, this design behaves like a band-pass filter in the topologic attention, supporting community-level diffusion over a learned relational manifold while maintaining structured contrast between clusters.
Figure 9: Learned Laplacian precision construction. The Laplacian matrix induces community-level smoothness and alternating cluster correlations.
@@
In the diagonally dominant formulation, a few high-confidence nodes can exert strong influence over their neighborhoods. Because each node’s self-term dominates its row in $J$, information primarily flows outward from confident nodes rather than through collective averaging. Each node is equipped with a learned confidence gate that dynamically adjusts its self-precision, regulating how much information it absorbs from its neighbors. As a result, highly confident nodes act as local leaders within the topologic attention, shaping how propagation unfolds across the graph. When node confidences are comparable, correlations weaken, leading to sparse yet focused connectivity patterns or minimal inter-node correlation. The more uniform edge spread observed in the adjacency view reflects this shift from community-level organization toward node-level importance. Overall, this design behaves as a low-pass filter centered around the most confident nodes.
Figure 10: Learned Diagonally dominant precision construction. The structure highlights sparse but concentrated correlations driven by high-confidence nodes.
@@
The pairwise normal matrix yields the most selective propagation pattern. Since each edge’s contribution depends on the mutual confidence of its connected nodes, information flows only where representations are jointly compatible. This formulation models each edge as the agreement both endpoints have in their connection, scaling similarity by their confidence ratio and emphasizing edges supported by reciprocal certainty. Consequently, propagation strength is highest along confident paths and suppressed in uncertain regions, requiring mutual trust for effective global communication. The resulting correlations form gradual gradients rather than sharp blocks, functioning as a high-pass filter in low-correlation areas and a low-pass filter in highly correlated areas. The corresponding adjacency maps reveal clearer separation between sparse and dense zones, indicating that low-degree nodes propagate information more readily than high-degree ones, where competition among edges dampens transmission. Overall, this design emphasizes strong, well-supported relationships while attenuating weak or inconsistent ones, producing structured selectivity in topologic attention.
Figure 11: Learned Pairwise normal precision construction. Information propagation is selective, reflecting mutual agreement between connected nodes.

### v2

@@ 3.3 Addressing the Unbounded Memory
Although GaBP guarantees convergence under walk-summability, the number of iterations required to reach convergence is unknown. This poses a challenge for training, as unrolling the iterative inference process would require storing, potentially, an unbounded number of intermediate messages for each layer. To address this, we adopt an implicit differentiation strategy that computes exact gradients at convergence without unrolling the iterations. At convergence, the graph linear transformation satisfies $J\mu=h$. Differentiating with respect to the model parameters $\theta$ yields
@@
This relation allows gradients to be computed directly at the fixed point using the same GaBP solver as in the forward pass, eliminating the need to store intermediate messages. This reduces memory complexity from $O(TE)$ to $O(E+N)$, where $T$ is the number of GaBP iterations, $E$ the number of edges, and $N$ the number of nodes, trading additional compute for substantial memory savings.
@@ 4 Experimental Setup
We evaluate the performance of learned and fixed topologies for graph linear transformation models across six benchmark datasets with cyclic graph structures under both homogeneous and heterogeneous learning settings. The Cora, Citeseer, and PubMed citation datasets serve as representative homogeneous benchmarks, while the Texas, Wisconsin, and Cornell datasets serve as representative heterogenous benchmarks. These six datasets each represent a single, static graph and are used for transductive node classification.
@@
We compare our method against foundational and representative models from dynamic neighborhood aggregation methods, along with standard local aggregation methods. Specifically, we include the Graph Convolutional Network (GCN) (Kipf and Welling, 2017), Graph Attention Network (GAT) (Veličković et al., 2018), and GraphSage (Hamilton et al., 2017) as representative local aggregation models. For methods that perform global or continuous neighborhood aggregation, we evaluate against the Structure-Aware Transformer (SAT) (Chen et al., 2022), an adapted Graphormer (Ying et al., 2021) designed for node classification rather than graph classification, as well as the Graph Neural Differential Equation (GDE) (Poli et al., 2019) and Continuous Graph Neural Network (CGNN) (Xhonneux et al., 2020).
@@
To ensure a fair comparison across models, we fix the hidden feature dimension to 64 across all methods. For models based on stacked aggregation layers, we use two aggregation layers. For head-based architectures, each layer employs eight heads in the first layer and a single head in the second layer. Each first layer head outputs eight features, while the second layer head outputs all 64 features. Continuous-depth models do not admit a discrete notion of layer depth; for these models, we parameterize the vector field in the same 64-dimensional hidden space. The goal of this setup is to control for model capacity so the differences primarily reflect the behavior of each aggregation mechanism rather than variations in network depth or parameter count.
@@
All models are trained using the Adam optimizer with weight decay $5\cdot 10^{-4}$, early stopping with a patience of 100 epochs and a dropout of $0.6$, except the fixed Laplacian which used a patience of 200 epochs. The learning rate for the graph linear transformation models and SAT models are $1\cdot 10^{-3}$ while the learning rate for the other models are $1\cdot 10^{-2}$. The tolerance for the GaBP was set to $1\cdot 10^{-6}$ and a maximum iteration count was set to 1000. To ensure stable convergence on graphs with loops, we apply message damping with a coefficient of $\lambda=0.5$. For more information on message damping, we provide the pseudocode algorithm for GaBP in the Appendix section A.2.
@@
For each dataset, we adopt random node splits preserving the original train/validation/test ratios but do not enforce class balance. This setup better reflects realistic scenarios where labeled nodes are unevenly distributed across classes, providing a more robust measure of generalization independent of split fragility (Shchur et al., 2019). Each experiment is repeated across 30 random runs, each with independently sampled splits. We report the mean $\pm$ standard deviation of test accuracy.
@@
For graph linear transformations, we ensure the graph is undirected and self loops are removed since the diagonal entries of the precision matrix $J$ already encode each node’s self-precision. For the other models, we ensure self loops.
@@ 5 Results
Table 1 reports test accuracy across the six benchmark datasets. Overall, the graph linear transformation models outperform or are on par to all baseline methods. However, learning topology is not always beneficial; the performance of the precision matrix depends on their respective inductive bias.
@@
Pairwise-normal constructions perform best on heterogeneous datasets, where selectively attending to informative edges mitigates the effects of heterophilic neighborhoods. In homophilic datasets, however, this selectivity restricts information flow, leading to reduced performance relative to less selective propagation mechanisms.
Laplacian precision matrices demonstrate the most stable performance across homophily levels, indicating that strong structural smoothing provides a robust prior for long-range information flow. Both fixed and learned Laplacian variants perform competitively on homophilic benchmarks while remaining effective in heterogeneous settings where local neighborhoods are unreliable, highlighting the benefits of community-level diffusion over aggressive edge filtering.
Diagonally dominant constructions exhibit a sharper trade-off between homophilic and heterogeneous graphs, with performance strongly dependent on whether node confidence is learned or fixed. Learned confidence mechanisms enable selective regulation of information flow, improving performance on heterogeneous datasets by suppressing misleading neighborhoods, but this same selectivity interferes with the uniform smoothing required for homophilic graphs, leading to degraded accuracy. In contrast, fixed diagonally dominant matrices align naturally with homophilic structure, where uniform confidence supports consistent feature diffusion but fail to correct erroneous propagation paths in heterogeneous graphs.
These behaviors are reflected in the correlation structures induced by each precision matrix (Appendix A.3): Laplacian constructions yield community-level correlation patterns, diagonally dominant matrices emphasize node centric influence, and pairwise-normal models induce highly selective pathways, collectively explaining the performance trends observed in Table 1.
5.1 Convergence Behavior
Table 2 reports the typical convergence iteration ranges of GaBP during training for each precision matrix. Because graph linear transformations rely on iterative inference, convergence behavior directly affects both computational efficiency and training dynamics.
Table 2: Typical GaBP iteration ranges required for forward and backward passes across precision matrix constructions. Values report observed ranges across datasets and training epochs.
Forward Pass Backward Pass
(L) Pairwise Normal 20 - 40 15 - 25
(L) Diagonally Dominant 25 - 100 20 - 40
(L) Laplacian 45 - 65 15 - 25
(F) Pairwise Normal 20 - 30 15 - 20
(F) Diagonally Dominant 70 - 200 20 - 25
(F) Laplacian >>1000 >>1000
Two clear trends emerge. First, except for the fixed Laplacian, the backward pass consistently converges faster than the forward pass across all constructions. This suggests that gradient propagation occurs over a smoother signal than the raw node observations used during forward inference.
Second, the fixed Laplacian is the only construction that does not converge within the allotted iteration budget, resulting in partially converged node mixing during both prediction and gradient computation. Despite this, the fixed Laplacian achieves the strongest overall performance. This indicates that full convergence may not be required for effective representation learning.
@@ 6 Conclusion
This work introduced Graph Linear Transformations, an approach for incorporating direct and indirect contexts in graphs. By interpreting graphs as walk-summable Gaussian graphical models, we are able to perform a linear transformation on the graph according to its topology. Across both homophilic and heterophilic benchmarks, graph linear transformations consistently outperforms all global aggregation methods, while matching or surpassing local neighborhood methods.
Two key challenges to broader adoption are improving the efficiency of graph linear transformations and expanding the class of Gaussian graphical models that can be represented on arbitrary graph structures. Since graph linear transformations rely on iterative methods, there is a clear path toward scalability through well-established acceleration techniques, such as multigrid methods (Briggs et al., 2000), preconditioning, and hierarchical solvers. Moreover, walk-summable graphs represent only a subset of valid Gaussian graphical models. Existing extensions of GaBP, such as those in (Johnson et al., 2009), demonstrate that inference can be performed on graphs that are not inherently walk-summable. We believe these directions offer promising avenues for future research and position graph linear transformations as a compelling alternative to both depth-based and dynamic aggregation methods.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Appendix
A.1 Implementation details of the Precision Matrices
The following paragraphs describe our implementations of each precision matrix. These paragraphs directly describe their corresponding precision matrix implementation in our github. However, we want to note that these are not the only way to implement the precision matrices, and we encourage the research into more well designed precision matrices.
Pairwise Normal Fixed
In standard Gaussian graphical models, the observation vector was typically used for defining the edges. Likewise for the fixed pairwise normal, we use the base input features to define the fixed pairwise normal matrix. The cosine similarity was the similarity function used to determine the strength of edges, $b_{ij}$, which was scaled to be between [0.99,0.99][-0.99,0.99]. The confidence for each node for each edge was set to 1 such that the pairwise constraint is always satisfied. Lastly, the matrix is symmetrically normalized.
Pairwise Normal Learned
In standard Gaussian graphical models, the observation vector was typically used for defining the edges. In this case, we use the learned observation vector to define our edge interactions. For the edge confidences, we let each node learn an unbounded self confidence for each edge pair. Thus, for an edge $i\to j$, we use a linear layer on the concatenation of the learned observation vector from both nodes: $a_{ij}=softplus([h_{i}||h_{j}])+1\cdot 10^{-8}$ and $c_{ij}=softplus([h_{j}||h_{i}])+1\cdot 10^{-8}$. To define the limits of the edge strength, we produce a scalar to scale the similarity function $scale=\epsilon\sqrt{a_{ij}c_{ij}}$, where $\epsilon$ is between [0, 1). We then scale the similarity function, cosine similarity in this case, using the scalar. Lastly, the matrix is symmetrically normalized.
Diagonally Dominant Fixed
For the fixed Diagonally Dominant matrix, we assume that each edge has a strength of 1. For the confidence of each node, we assume it is 1 plus the degree of the node.
Diagonally Dominant Learned
For the learned Diagonally Dominant matrix, we reuse the learned observation vector for the similarity vector. For the similarity function, we use the cosine similarity. For the learned self confidence, we take the original node embeddings $X$ and use them to produce node unbounded specific confidences: $J_{ii}=\text{Softplus}(XW)$. We then symmetric normalize the matrix.
Laplacian Fixed
For the fixed Laplacian, we take the normal adjacency matrix and have the diagonal be $1\cdot 10^{-8}$ + the degree of the node. We then symmetrically normalize the matrix and scale the off diagonals by 0.99 so the eigenvalues are between [-0.99, 0.99].
Laplacian Learned
For the learned Laplacian, we use a separate observation vector and similarity vector. Edge weights are defined using a Gaussian kernel on the learned similarity embeddings, producing nonnegative edge strengths. These weights are degree-normalized to form a Laplacian-style precision matrix, where off-diagonal entries encode normalized negative interactions. To ensure walk-summability, the off-diagonal terms are scaled to lie strictly within (1,1)(-1,1), and the diagonal entries consist of a fixed spectral floor plus a small learned node-specific confidence. The learned diagonal contribution is bounded to prevent it from dominating the Laplacian structure. Finally, the matrix is symmetrically normalized and then scaled by a learned global sigmoid factor, following the normalization principles of (Kipf and Welling, 2017).
A.2 Pseudocode for Gaussian Belief Propagation
During GaBP, message damping can be beneficial for improving convergence stability. Damping replaces each message update with a convex combination of its previous value and the newly computed value, effectively acting as a moving average over iterations. This is particularly helpful on frustrated graphs, where competing edge interactions cannot be simultaneously satisfied and undamped message updates may oscillate.
@@ A.2 Pseudocode for Gaussian Belief Propagation
8:  $\Delta\pi_{i\to j}=\pi^{\text{new}}_{i\to j}-\pi_{i\to j}$
9:  $\Delta\eta_{i\to j}=\eta^{\text{new}}_{i\to j}-\eta_{i\to j}$
10:  $\pi_{i\to j}\leftarrow(1-\lambda)\,\pi_{i\to j}+\lambda\,\pi^{\text{new}}_{i\to j}$
11:  $\eta_{i\to j}\leftarrow(1-\lambda)\,\eta_{i\to j}+\lambda\,\eta^{\text{new}}_{i\to j}$
12:  $\Delta=\max_{(i,j)\in E}\left(\max\left(|\Delta\pi_{i\to j}|,\;|\Delta\eta_{i\to j}|\right)\right)$
13:end while
14:Output: $\mu_{i}=\eta_{i}/\pi_{i}$
Algorithm 1 provides the pseudocode for GaBP with message damping and serves as a reference for the equations presented in the background section.
@@ A.3 Analysis of the Biases of Matrix Constructions
To understand how each precision matrix construction shapes the topology, we visualize the correlation matrix (Figures 47) on the Wisconsin dataset. For the learned precision matrices, we visualize the first-layer. These correlation matrices are computed from the inverse precision matrix and describe how strongly two nodes would co-vary under the learned topology in the absence of any observation vector $h$. They therefore characterize the influence structure implied by $J$, meaning the potential pathways and relative strengths along which GaBP is capable of transmitting information, regardless of the actual evidence supplied during inference. Red values indicate positive correlations and blue values indicate negative correlations, with intensity proportional to magnitude. Nodes are ordered by the second eigenvector to highlight community structure, and the accompanying adjacency matrices (bottom panels for Figure  4 and right panels for Figures 5-7), ordered in the same way, show the corresponding learned connectivity patterns.
@@
Fixed Constructions.
The fixed precision matrices show that the Laplacian and diagonally dominant are more homogenous in their representation, while the fixed pairwise normal is quite selective in its neighboring representations.
@@
The Laplacian-based precision matrix produces a pronounced block structure resembling a plaid pattern of alternating positive and negative correlations. Nodes form compact clusters with internal correlations, while inter-cluster relationships often alternate in sign, indicating that the model captures both homophilic and heterophilic dependencies. The dense regions along the diagonal of the reordered adjacency matrix confirm the emergence of localized communities. Because the Laplacian construction constrains the eigenvalues of the precision matrix to lie within (1,1)(-1,1), each node contributes with comparable influence to the overall propagation. Most of the learning therefore occurs through topological deformation, where the model folds the underlying space before blending node representations. As a result, this design behaves like a band-pass filter, supporting community-level diffusion over a learned relational manifold while maintaining structured contrast between clusters. This can be seen with the different block structures present in the adjacency matrices.
Figure 5: Learned Laplacian precision construction.
@@
In the diagonally dominant formulation, a few high-confidence nodes can exert strong influence over their neighborhoods. Because each node’s self-term dominates its row in $J$, information primarily flows outward from confident nodes rather than through collective averaging. Each node is equipped with a learned confidence gate that dynamically adjusts its self-precision, regulating how much information it absorbs from its neighbors. As a result, highly confident nodes act as local leaders. When node confidences are comparable, correlations weaken, leading to sparse yet focused connectivity patterns or minimal inter-node correlation. The more uniform edge spread, observed in the adjacency matrix, reflects this shift from community-level organization toward node-level importance. Overall, this design behaves as a low-pass filter centered around the most confident nodes.
Figure 6: Learned Diagonally dominant precision construction.
@@
The pairwise normal matrix yields the most selective propagation pattern. Since each edge’s contribution depends on the mutual confidence of its connected nodes, information flows only where representations are jointly compatible. This formulation models each edge as the agreement both endpoints have in their connection, scaling similarity by their confidence ratio and emphasizing edges supported by reciprocal certainty. Consequently, propagation strength is highest along confident paths and suppressed in uncertain regions, requiring mutual trust for effective global communication. The resulting correlations form gradual gradients rather than sharp blocks, functioning as a high-pass filter in low-correlation areas and a low-pass filter in highly correlated areas. The corresponding adjacency matrix reveal clearer separation between sparse and dense zones, indicating that low-degree nodes propagate information more readily than high-degree ones, where competition among edges dampens transmission.
Figure 7: Learned Pairwise normal precision construction.